OpenStudy (anonymous):

Help please? Suppose you have a normally distributed set of data pertaining to a standardized test. The mean score is 500 and the standard deviation is 100. What is the z-score of 450 point score? 1.5 1 0.5 -0.5

OpenStudy (kropot72):

Use the formula \[\large z=\frac{X-\mu}{\sigma}\] where X is the point score, mu is the population mean, and sigma is the standard deviation.

OpenStudy (kropot72):

Plugging in the given values: \[\large z=\frac{450-500}{100}=you\ can\ calculate\]

OpenStudy (anonymous):

thank you!

OpenStudy (kropot72):

You're welcome :)